### v1

@@ 4.2 System Operation
Summary rules. All mint transactions are summarized as follows: a list of liquidity positions will be produced with each position consisting of a tuple containing: the position identifier, the public key of the owner, total amount of liquidity provided (or net change) for each token type under this position, and total amount of accrued fees. Note that the payin of the LP is also updated when summarizing mint transactions; all provided liquidity token amounts are deducted from their deposit and reflected on the payin structure.
Burns. Burn transactions consist of a partial or complete liquidity withdrawal from a position (the latter will result in deleting the position). In order to execute a burn transaction, an LP broadcasts their burn transaction containing the following inputs: the position ID, the tick price limits and the desired amounts of tokens $A$ and $B$ to be burned. The algorithm then calculates the amount of liquidity that can be burnt from the position, given those inputs, and then returns the calculated amounts of tokens $A$ and $B$.
@@
Processing. Processing a burn transaction boils down to determining if the issuer LP owns the position, then calculating the amount of liquidity this LP owns in a share, and converting that amount of liquidity into an amount of both tokens. This would lead to an update to the position range (upper and lower price ticks), or a deletion of the position if all its associated liquidity is withdrawn. If a deleted position has fees owed to it, the owner LP will receive these fees as part of her total payout computed at the end of the epoch, i.e., will be added to their sidechain deposit.
@@
Summary rules. Burns are summarized as part of summarizing mint transactions detailed above. Burns adjust the net changes of the pool liquidity balance, i.e., they decrement this balance. Any fully withdrawn position, i.e., has zero remaining liquidity balance, will be removed from the $\mathsf{TokenBank}$ state. The payin and payout for each LP will be computed as done for mints above. That is, provided liquidity will be deducted from the corresponding deposit that covered the mint transactions, and withdrawn liquidity will be added to sidechain deposit balance to be reflected on the payout.
@@
Collects. Collect transactions allow LPs to collected the fees earned by their liquidity positions. An LP broadcasts a collect transaction containing the identifier of the position and the amount of fees to collect.
@@
Processing. This includes determining if the issuer LP owns the position, and checking if the amount they want to collect can be covered by their fee balance. If all is fine, the LP sidechain deposit of the issuer LP is updated to reflect the amount of collected fees, and the fee balance for that position is adjusted accordingly.
@@
Summary rules. Summarizing collect transactions is also part of summarizing mints/burns and the LP payout structure. That is, fee balance of the referenced position is decreased based on the collected amount, and the payout to the issuer LP is computed based on the remaining sidechain deposit balance (to which the collected fee amount has been added).
@@
Flashes. Flash transactions allow users to request short-term loans within the duration of one mainchain block. Flashes are the only transaction type which $\mathsf{ammBoost}$ does not offload to the sidechain; the delay in paying out the actual tokens (which happens at the end of an epoch) limits the intended use of flash loans that span a very short period. As such, in $\mathsf{ammBoost}$, flash transactions happen on the mainchain as in the original AMM architecture. Since flash loans take place in a singular block, they do not impact the pool balances; the amount of loaned tokens should be returned within one block period or the loan will be inverted. As a result, they do not invalidate any of the transactions processed on the sidechain based off the balance snapshot taken at the beginning of an epoch.
@@
Syncing $\mathsf{TokenBank}$. After producing a summary-block containing all the summaries detailed earlier, the sidechain committee leader calls the $\mathsf{Sync}$ function in $\mathsf{TokenBank}$ that resembles a sync-transaction submission. The inputs to this function call include: the list of payouts and payins for all clients and LPs, the list of liquidity positions with their updated information (i.e., balances of each token type, price range, and fees), and the updated pool balances.
@@
Authentication. $\mathsf{TokenBank}$ must ensure that the $\mathsf{Sync}$ function invocation is issued by the rightful sidechain committee of the epoch. We use a modified idea of quorum certificates (QC) [52, 65] combined with threshold signatures. In detail, to authenticate the $\mathsf{Sync}$ call for epoch $e+1$, the election of committee $e+1$ must happen during epoch $e$. Then, this committee runs a distributed key generation (DKG) [34] to generate a public verification key $\mathsf{vk}_{c}$ for the committee and shares of the signing key (one share per committee member) with a threshold of $22f+$ 2 (so a committee size is $22f+$ 2 and $f$ is the maximum number of faulty nodes as in PBFT protocols). This committee initiates an agreement on $\mathsf{vk}_{c}$, and then sends agreement output to the committee $e$ along with their proofs of election of each member participated in the agreement. [^4] In our implementation, this election proof is basically the output of the verifiable random function (VRF) used in the election mechanism. Committee $e$ verifies the election proofs, and then verifies that there is an agreement on $\mathsf{vk}_{c}$. If everything is correct, committee $e$ records $\mathsf{vk}_{c}$ on $\mathsf{TokenBank}$ by adding that to the $\mathsf{Sync}$ function call inputs submitted at the end of epoch $e$. During epoch $e+1$, committee $e+1$ runs an agreement over the $\mathsf{Sync}$ function call inputs (which are the summaries used for syncing as detailed before) and sign using there signing key shares, which result in one signature over these inputs. The leader them invokes $\mathsf{Sync}$ with the inputs and this signature. In turn, $\mathsf{TokenBank}$ verifies the signature using the recorded $\mathsf{vk}_{c}$ before accepting the summaries. By the security of the threshold signature scheme, this signature will be valid only if at least the required threshold of the committee has signed.
@@
Processing. If successfully verified, $\mathsf{TokenBank}$ processes the $\mathsf{Sync}$ function call by updating the list of positions based on the summaries, i.e., delete fully withdrawn positions, create new positions, or adjust the parameters of existing ones. Lastly, it updates the pool balance based on the reported values. Then it dispenses the payouts to the referenced clients/LPs, deducts payins from their deposits, and refunds the residual of the deposits (if any) to their owners. Note that since we allow users to use the newly claimed tokens (tracked using the sidechain deposits) in transactions during an epoch, if a total payin of a particular user exceeds the mainchain deposit, the rest is deducted from the total payouts this user will get.
@@
Handling interruptions. We identify the scenarios that can lead to operation interruption in $\mathsf{ammBoost}$ and how to recover from them. Recall that the sidechain committees use a PBFT-based consensus that assumes up to $f$ of the elected miners can be malicious. [^5] This is valid under a committee size that guarantees satisfying this condition with overwhelming probability, where we adopt the committee size analysis from [54]. Thus, interruptions that could happen result from having a malicious or unresponsive leader. This leader may either propose an invalid meta/summary-blocks or invalid function call to $\mathsf{Sync}$, or not initiate the agreement in the first place. Another interruption could result from rollbacks on the mainchain. That is, when the mainchain miners switch their canonical chain to track the one satisfying the fork resolution system of the blockchain system (i.e. the longest blockchain, or the heaviest one), causing the most recent blocks to be abandoned. This is an issue if the abandoned blocks contain $\mathsf{Sync}$ transactions.
Detection and recovery from these interruptions are done as proposed in chainBoost (which we review briefly here while full details can be found in [54]). A leader that proposes an invalid block or $\mathsf{Sync}$ call inputs, can be easily detected by the committee by verifying the proposed blocks and the syncing inputs. Once detected, the view-change technique [40] is used to elect a new leader. In case of unresponsive leader, there is a time out $\tau$ so that if no agreement is initiated within that time, this triggers a leader change. The outcome of the view-change operation is recorded on the meta-blocks and is included in the summary-block for accountability. As for a leader that proposes invalid $\mathsf{Sync}$ inputs, a leader-change will not help since this happen at the end of an epoch when it is time for the new committee to take over. We handle this case, and the rollback interruption, using the mass-syncing technique. The new committee issues a $\mathsf{Sync}$ call covering the summaries they produced in their epoch and those produced earlier in the impacted epochs.
@@ 4.3 Security
Since $\mathsf{ammBoost}$ delegates the processing of the AMM transactions to the sidechain, and introduces pruning and state synchronization, we show that under this new architecture the security and correct operation (i.e., safety and liveness) of the underlying AMM are preserved. In Appendix 0.A, we prove the following theorem:
@@ 5 Implementation
To assess the performance gains that $\mathsf{ammBoost}$ provides for AMMs, we implement a proof-of-concept and conduct a varied set of performance evaluations. We chose a Uniswap-inspired use case to represent the underlying AMM. This section discusses the implementation details, while the experimental setup and results are discussed in the next section.
@@
Sidechain Implementation. For the sidechain, we use the chainBoost implementation from [2]. This code uses cryptographic sortition-inspired election mechanism [47] for the committee election, and this committee runs a BLS collective signing (CoSi)-based PBFT-based consensus algorithm [5]. We add the modifications needed in $\mathsf{ammBoost}$; our sidechain implementation deploys its own miners, and adopts the summary rules defined in Section 4. We also modify the syncing process to be an invocation to the $\mathsf{Sync}$ function in $\mathsf{TokenBank}$ authenticated using the threshold signature-based quorum certificate approach discussed earlier. We use a simplified version of the golang BLS library [62], and a pre-generated key to sign the $\mathsf{Sync}$ transaction. Furthermore, our sidechain implements two extra functions to aid in performing the AMM functionality:
@@
  • $\mathsf{CreateTx}_{\mathsf{Sync}}$: creates the inputs of the $\mathsf{Sync}$ call using the information found in a summary-block. This is called by the sidechain committee leader at the end of the epoch.
  • $\mathsf{SnapshotBank}:$ retrieves users’ deposits at the beginning of an epoch.
@@
Mainchain details. For the mainchain, we utilize the Ethereum Sepolia testnet [18] using the hardhat development environment [12]. We implemented the $\mathsf{TokenBank}$ smart contract in Solidity [19] and deployed it on Sepolia. $\mathsf{TokenBank}$ handles users’ deposit, storage of tokens, and the syncing and payout processes. In our implementation, the interfacing between the sidechain miners and the mainchain is handled through functionality provided by the Go-Ethereum project [11]. To allow $\mathsf{TokenBank}$ to authenticate the $\mathsf{Sync}$ function call, we implement BLS signature verification in solidity, where we use the implementation of the 256-bit Barreto-Naehrig (BN256) curve operations defined in the Ethereum precompiles [59, 39]. We implement our hash-to-point functionality as the scalar multiplication of a Keccak256 hash of the $\mathsf{Sync}$ entries and the generator of the $G_{2}$ curve of BN256.
@@
Use case: Uniswap-inspired AMM. We implement swaps, mints, burns, and collects using the same logic as in Uniswap (Appendix 0.B). We do not implement flashes since they make up a small portion of the traffic so they will not impact the performance gains we report. For simplicity, our use case implementation manages a single pool. We deployed two standard ERC20 contracts to provide the token pair traded in this pool and used in both the ammBoost and baseline experiments. Naturally, to test Uniswap V3 in an isolated environment, we used the UniswapV3Factory contract to deploy a new liquidity pool which hosts the two ERC20 tokens. In order to test against the baseline implementation of Uniswap V3, we wrote and deployed a smart contract to interface with the various Uniswap contracts as detailed in the Uniswap documentation [25]. This smart contract, from here on out referred to as the interface contract, acts as our point of contact for the Uniswap smart contracts. The interface contract routes swaps to the swapRouter contract and mints/burns/collects to the NFPM. Additionally it receives and keeps track of all of the NFT liquidity positions (through the ERC721Reciever interface) created by the users in our experiments.
@@ 6 Performance Evaluation
Experiment setup. We deploy our system on a computing cluster composed of 8 hypervisors, each running a 12-Core, 130 GiB RAM, VM, connected with 1 Gbps network link. This setup is capable of running around 8000 sidechain miners. Unless stated otherwise, an experiment length is 11 epochs with each epoch consists of 10 mainchain rounds (which corresponds to 30 sidechain rounds per epoch). Our default meta-block size is 1 MB and a sidechain committee contains 500 miners. We deploy 100 users to participate in the AMM, generating traffic that arrives at a constant rate of $\rho=\lceil\frac{V_{D}\times b_{t}}{3600\times 24}\rceil$ where $V_{D}=25\times 10^{6}$ is the chosen daily volume of transactions. We benchmark our system against a baseline, which is a deployment of Uniswap V3 on Sepolia, as mentioned earlier. We test the impact of multiple parameters on $\mathsf{ammBoost}$’s performance, including various traffic amounts and distributions, block size, sidechain round duration, and number of sidechain rounds per epoch. In reporting the results, we measure the following metrics:
@@
    Throughput: The number of transactions processed per mainchain round.
@@
    Transaction latency: The delay between a transaction submission and its appearance in a meta-block. To obtain an accurate representation of this metric, and thus process a comparable amount of traffic, we empty the transaction queues after the end of each run.
@@
    Payout latency: The delay between the submission of a transaction and the completion of $\mathsf{TokenBank}$ syncing using the summary-block of the epoch in which this transaction has been published on the sidechain. We measure this metric by reporting the sum of the transaction latency (previous metric), the time needed to issue a $\mathsf{Sync}$ call for the epoch during which a transaction appeared in a meta-block, and the time needed to process the transaction encapsulating the $\mathsf{Sync}$ call.
@@
    Gas cost: The average number of gas units paid to process critical transactions (such as user deposits and the $\mathsf{Sync}$ call).
@@
Scalability. In this experiment, we test the scalability of a single $\mathsf{ammBoost}$ pool to understand its behavior under heavy traffic. We follow the same traffic distribution as in Uniswap and vary the daily volume $V_{D}\in\{50K,500K,5M,25M\}$. We record the impact on the throughput, as well as transaction and payout latency. Our results are shown in Table 1.
Table 1: Scalability of $\mathsf{ammBoost}$.
Daily volume 5K 50K 500K 25M
Throughput (tx/s) 0.75 6.0 58.0 242.35
Average sidechain latency (s) 0.075 0.075 0.075 128.29
Average payout latency (s) 75.53 75.53 75.53 204.54
Table 2: Impact of different sidechain block sizes.
BlockSize (MB) 0.5 1 1.5 2
Throughput (tx/s) 121.71 241.61 362.52 483.43
Average sidechain latency (s) 2487.35 913.20 389.89 128.03
Average payout latency (s) 2562.90 988.89 466.01 203.64
Throughput wise, we record a low throughput of 0.75 tx/s up to 78 tx/s when the daily volume is between 50K and 5M transactions (roughly equivalent to 1x-100x Uniswap’s daily volume). This is mainly due to the mainchain blocks not being full as this workload is way below the capacity that our system can handle. The maximum throughput we can achieve is 242.35 tx/s when then traffic is 500x Uniswap’s daily volume.
Latency wise, we achieve a quasi-instant when the daily volume is between 50K and 5M (transactions that arrive at the beginning of the round get processed within the same round, while the residual amount of latency is due to transactions generated at the end of the epoch and processed in the next epoch). This leads to payouts being processed within one epoch (including the time needed to confirm the transaction encapsulating the $\mathsf{Sync}$ function call on the mainchain). Transaction congestion happens when the daily volume is 25M, resulting in higher average transaction and payout latency as the table shows.
Impact of block size. In this experiment, we test the impact of the sidechain block size. As this experiment is primarily geared towards finding an optimal block size for our system, we compare different deployments of $\mathsf{ammBoost}$ with different block sizes against Uniswap’s Sepolia deployment. We run the protocol with the following block sizes {$0.5,1,1.5,2$} MB. We measure the impact on throughput and transaction and payout latency with the goal of identifying the block size that maximizes throughput while minimizing latency. For this experiment, we increase the daily volume to 50M transactions. Our results are in Table 2.
As shown, throughput increases as the block size increases since more transactions can be packed in a block. Increasing the block size also improves latency, as packing more transactions in a block reduces the queue congestion. However, larger block sizes mean a larger propagation delay that could be problematic for short sidechain round duration. Thus, system designers should be careful when choosing an optimal sidechain block size, balancing between the block size needed for the volume of transactions they are facing while capturing the intricacies of large network transfers.
Impact of sidechain round duration. In this experiment, we study the impact of the sidechain round duration. An ideal round duration should allow for consensus to conclude while maximizing throughput and minimizing latency. Inspired by Polygon [29] (which has a block time of 2 seconds and Algorand [13] which has a block time of around 3.6 seconds), we test the following round duration values: 4, 6, 9, and 12 seconds, and report the performance metrics as before (Table 3).
Table 3: Impact of sidechain round duration.
Sidechain round duration (s) 4 6 9 12
Throughput (tx/s) 241.61 161.33 108.06 81.05
Average sidechain latency (s) 128.29 520.65 1105.81 1688.51
Payout latency (s) 204.54 628.23 1260.17 1889.47
Table 4: Impact of number of sidechain (sc) rounds per epoch.
Epoch Length (sc rounds) 5 10 20 30 60 96
Throughput (tx/s) 199.98 224.94 237.82 241.61 246.16 247.68
Sidechain latency (s) 291.98 186.59 142.04 128.29 114.96 110.03
Payout latency (s) 318.04 222.47 197.87 204.54 255.08 318.57
Table 5: Impact of traffic distribution.
Swap % 60 80
(Mint %, Burn %, Collect %) (10,20,10) (10,20,10) (10,10,20) (10,5,5) (5,10,5) (5,5,10)
Throughput (tx/s) 254.04 251.59 251.22 246.59 245.40 245.24
Sidechain latency (s) 88.72 96.20 97.36 111.70 118.89 116.89
Payout latency(s) 165.40 172.86 173.99 187.82 195.01 191.82
Max sidechain growth (B) 31831 31831 31831 31831 31831 31831
Throughput wise, we observe that as the block time increases, throughput decreases. This is due to processing the same amount of transactions while increasing the time needed to produce a block. Increasing block time also increases latency as it takes longer to have transactions published in a block. To choose the optimal block time, system designers need to take into consideration the time required for the sidechain consensus and network propagation delays, while aiming to generate new blocks as fast as possible.
Impact of the number of sidechain rounds per epoch. Here, we test the performance impact of the number of sidechain rounds within an epoch. The goal is to find an epoch that maximizes throughput, minimizes transaction latency and payout delay, based on the optimal sidechain round duration from the experiment above. Thus, we pick our epoch to have $\{5,10,20,30,60,96\}$ sidechain rounds, and report the performance metrics (Table 4)
Having very short epochs negatively affects throughput and the sidechain latency. That is, frequent summary-blocks harm performance since this leads to a larger number of $\mathsf{Sync}$ calls that are costly. At the same time, fewer transactions are processed within epoch, thus affecting both latency and throughput. Longer epoch duration reduces the sidechain latency and increases throughput. However, this affects the payout latency adversely since $\mathsf{Sync}$ calls now are much fewer and spaced out (so users have to wait longer, as the epoch itself is longer, to obtain their actual token payouts). Also, this means that these users have to put larger deposits to cover their (long) epoch activities, which could be undesirable. Based on our results, we achieve the best payout latency when the epoch lasts for 20 sidechain rounds, which is equivalent to 80 sec.
Impact of traffic distribution. In this experiment, we evaluate different traffic distributions as follows (where we keep the swap operations dominant and all the number are percentages): $(s,m,b,c)\in\{(60,20,10,10)$, $(60,10,20,10),$ $(60,10,10,20),$ $(80,10,5,5),$ $(80,5,10,5),$ (80,5,5,10)}(80,5,5,10)\}( 80 , 5 , 5 , 10 ) }, where $(s,m,b,c)$ stand for swaps, mints, burns, and collects, respectively. Our results can be found in Table 5.

### v3

@@ 4.2 System Operation—Transaction Processing
Summary rules. All mint transactions are summarized as a list of liquidity positions with each position consisting of a tuple containing: the position identifier, the public key of the owner, the total amount of liquidity provided (or net change) for each token type under this position, and total amount of accrued fees. Note that the payin/payout of the LP is also updated when summarizing mint transactions; all provided liquidity token amounts are deducted from their deposits as shown in Figure 5.
Burns. A burn transaction allows a partial or complete liquidity withdrawal of a position. It is issued by an LP and contains: the position ID, the tick price limits, and the desired amounts of tokens $A$ and $B$ to be burned, and sent to the sidechain.
@@
Processing. Processing a burn transaction boils down to determining if the issuer LP owns the position, then calculating the amount of liquidity this LP owns in a share, and converting that amount of liquidity into an amount of both tokens managed by the pool (using the original logic of the AMM). This would lead to updating the position range (upper and lower price ticks), or a deletion of the position if all its associated liquidity is withdrawn. If a deleted position has fees owed to it, the owner LP will receive these fees as part of her total payout computed at the end of the epoch, i.e., will be added to their deposit balance.
@@
Summary rules. Burns are summarized as part of summarizing mint transactions detailed above. Burns adjust the net changes of the pool liquidity balance, i.e., they decrement this balance. Any fully withdrawn position will be removed from the $\mathsf{TokenBank}$ state. The withdrawn liquidity will be added to the LP’s deposit balance to be reflected on the payout.
@@
Collects. Collects allow LPs to collect the fees earned by their liquidity positions. An LP broadcasts a collect transaction containing the identifier of the position and fee amount to be collected.
@@
Processing. This includes determining if the issuer LP owns the position, and checking if the amount they want to collect can be covered by their fee balance. If all is fine, the issuer LP’s deposit is updated to reflect the amount of collected fees, and the fee balance for that position is adjusted accordingly.
@@
Summary rules. Summarizing collect transactions is also part of summarizing mints/burns and the LP payout structure. That is, fee balance of the referenced position is decreased based on the collected amount, and the payout to the issuer LP is computed based on their updated deposit balance (to which the collected fee amount has been added).
@@
Flashes. Flash transactions allow users to request short-term loans within the duration of one mainchain block. These are the only transaction type that $\mathsf{ammBoost}$ does not offload to the sidechain; the delay in paying out the actual tokens (which happens at the end of an epoch) limits the intended use of flash loans that span a very short period. As such, in $\mathsf{ammBoost}$, flash transactions happen on the mainchain as in the original AMM architecture. Since flash loans take place in a singular block, they do not impact the pool balances; the amount of loaned tokens should be returned within one block period or the loan will be inverted. As a result, they do not invalidate any of the transactions processed on the sidechain based off the balance snapshot taken at the beginning of an epoch.
Remark 1
In terms of user experience, clients and LPs should be connected to both the mainchain and the sidechain, and their wallets should issue transactions to the destination chain based on the transaction type. For example, deposits should go the mainchain, while swaps/mints/burns/collects should go to the sidechain. Another difference is related to receiving the actual tokens, which are delayed until the end of the epoch in $\mathsf{ammBoost}$. Since a user can use these immediately for trades within an epoch, the delayed payout has no impact. However, if a user wants to use these tokens on a different AMM, or wants to trade these with other token types to participate in another pool managing different token pairs, then they have to wait until the epoch. Still, overall, the delay is only one epoch.
4.3 System Operation—Chain Management
In this section, we discuss the syncing process, sidechain pruning, and how $\mathsf{ammBoost}$ recovers from interruptions.
@@ 4.3 System Operation—Chain Management
Syncing $\mathsf{TokenBank}$. The sidechain committee leader, after producing a summary-block containing all the summaries detailed earlier, calls the $\mathsf{Sync}$ function in $\mathsf{TokenBank}$ that resembles a sync-transaction submission. The inputs to this function call include: the list of payouts for all clients and LPs, and the list of liquidity positions with their updated information.
@@
Authentication. $\mathsf{TokenBank}$ must ensure that the the $\mathsf{Sync}$ function invocation is issued by the rightful sidechain committee. We use a modified idea of quorum certificates (QC) [54, 72] combined with threshold signatures. In detail, to authenticate the $\mathsf{Sync}$ call for epoch $e+1$, the election of committee $e+1$ must happen during epoch $e$. Then, this committee runs a distributed key generation (DKG) [35] to generate a public verification key $\mathsf{vk}_{c}$ for the committee and secret shares of the signing key (one share per member) with a threshold of $22f+$ 2.999A committee size is $23f+$ 2 and $f$ is the maximum number of faulty nodes as in PBFT protocols, $22f+$ 2 votes are needed to reach an agreement. This committee initiates an agreement on $\mathsf{vk}_{c}$, and then sends the agreement output to committee $e$ along with proofs of election of each member who participated in the agreement. [^10] In our implementation, this election proof is the output of the verifiable random function (VRF) used in the election mechanism. Committee $e$ verifies the election proofs, and then verifies that there is an agreement on $\mathsf{vk}_{c}$. If everything is correct, committee $e$ records $\mathsf{vk}_{c}$ on $\mathsf{TokenBank}$ by adding that to the $\mathsf{Sync}$ function call inputs they submit at the end of epoch $e$. During epoch $e+1$, committee $e+1$ runs an agreement over the $\mathsf{Sync}$ function call inputs and signs using their signing key shares, which result in one signature over these inputs. The leader then invokes $\mathsf{Sync}$ with the inputs and this signature. In turn, $\mathsf{TokenBank}$ verifies the signature using the recorded $\mathsf{vk}_{c}$ before accepting the summaries. By the security of the threshold signature scheme, this signature will be valid only if at least $22f+$ 2 committee members has signed.
@@
Processing. If successfully verified, $\mathsf{TokenBank}$ processes the $\mathsf{Sync}$ function call by updating the list of positions based on the summaries, i.e., delete fully withdrawn positions, create new positions, or adjust existing ones, Then, it updates the pool balance based on the reported payouts and updated position list. Lastly, it dispenses the payouts to the referenced clients/LPs.
@@
Handling interruptions. We identify the scenarios that can lead to operation interruption in $\mathsf{ammBoost}$ and how to recover from them. Recall that the sidechain committees use a PBFT-based consensus that assumes up to $f$ of the elected miners can be malicious. [^11] This is valid under a committee size that guarantees satisfying this condition with overwhelming probability, where we adopt the committee size analysis from [57]. Thus, interruptions that could happen result from having a malicious or unresponsive leader. This leader may either propose an invalid meta/summary-blocks or invalid function call to $\mathsf{Sync}$, or not initiate the agreement in the first place. Another interruption could result from rollbacks on the mainchain. That is, when the mainchain miners switch their canonical chain to the one satisfying a particular fork resolution criteria (i.e. the longest branch, or the heaviest one), causing the most recent blocks to be abandoned. This is an issue if the abandoned blocks contain $\mathsf{Sync}$ transactions.
Detection and recovery from these interruptions are done as in chainBoost [57], which we review briefly here. A leader that proposes an invalid block or $\mathsf{Sync}$ call can be easily detected by the committee when verifying this proposal. Once detected, the view-change technique [41] is used to elect a new leader. In the case of an unresponsive leader, if no agreement is initiated within a timeout period, a leader change is triggered. As for a leader that proposes invalid $\mathsf{Sync}$ inputs, a leader-change will not help since this happens at the end of an epoch when it is time for the new committee to take over. Thus, this case, and the rollback interruption, are addressed using the mass-syncing technique. The new committee issues a $\mathsf{Sync}$ call covering the summaries they produced in their epoch and those produced earlier in the impacted epochs.
Remark 2
A sidechain operates as a regular blockchain, thus any transactions that have not been processed in an epoch will be carried over to the epoch after. All sidechain miners receive transactions destined to the sidechain, but only the elected committee mines meta and summary blocks. Thus, when a new meta-block is mined, the committee and all other sidechain miners remove all published transactions in that block from their queues.
@@ 4.4 Security
Since $\mathsf{ammBoost}$ delegates the processing of the AMM transactions to the sidechain, and introduces pruning and state synchronization, we show that under this new architecture, the security and correct operation (i.e., safety and liveness) of the underlying AMM are preserved. In Appendix 0.A, we prove the following theorem:
@@ 5 Implementation
To assess the performance gains that $\mathsf{ammBoost}$ provides for AMMs, we implement a proof-of-concept and conduct various experiments. [^12] We will open source our implementation. We chose a Uniswap-inspired use case to represent the underlying AMM. This section discusses the implementation details, while the experimental setup and results are discussed in the next section.
@@
Sidechain implementation. For the sidechain, we use the chainBoost implementation from [3]. This code uses cryptographic sortition-inspired election mechanism [48] for the committee election, and this committee runs a BLS collective signing (CoSi)-based PBFT-based consensus algorithm [7]. We add the modifications needed for $\mathsf{ammBoost}$; our sidechain has its own miners, and adopts the summary rules defined in Section 4. We also modify the syncing process to be an invocation to the $\mathsf{Sync}$ function in $\mathsf{TokenBank}$ authenticated using the threshold signature-based quorum certificate discussed earlier. We use a simplified version of the golang BLS library [68], and a pre-generated key to sign the $\mathsf{Sync}$ transaction. Furthermore, our sidechain implements two extra functions to aid in performing the AMM functionality:
@@
  • $\mathsf{CreateTx}_{\mathsf{Sync}}$: creates the $\mathsf{Sync}$ call inputs based on the summary-block, including the payouts, the updated liquidity positions, and the pool liquidity balance. This is called by the sidechain committee leader at the end of each epoch.
[…]
@@
Mainchain details. For the mainchain, we utilize the Ethereum Sepolia testnet [19] using the hardhat development environment [14]. We implemented $\mathsf{TokenBank}$ in Solidity [20] and deployed it on Sepolia. In our implementation, the interfacing between the sidechain miners and the mainchain is handled through functionality provided by the Go-Ethereum project [13]. To allow $\mathsf{TokenBank}$ to authenticate the $\mathsf{Sync}$ function call, we implement BLS signature verification in solidity, where we use the 256-bit Barreto-Naehrig (BN256) curve operations defined in the Ethereum precompiles [64, 40]. We implement our hash-to-point functionality as the scalar multiplication of a Keccak256 hash of the $\mathsf{Sync}$ entries and the generator of the $G_{2}$ curve of BN256.
@@
Use case: Uniswap-inspired AMM. We implement swaps, mints, burns, and collects using the same logic as in Uniswap (Appendix 0.B). We do not implement flashes since they represent a very small portion of the traffic, and thus will not impact the performance gains we report. For simplicity, our use case implementation manages a single pool. We deployed two standard ERC20 contracts to provide the token pair traded in this pool and used in both the ammBoost and baseline experiments. Naturally, to test Uniswap V3 in an isolated environment, we use the UniswapV3Factory contract to deploy a new liquidity pool which hosts the two ERC20 tokens. In order to test against the baseline implementation of Uniswap V3, we wrote and deployed a smart contract to interface with the various Uniswap contracts as detailed in the Uniswap documentation [26]. This interface contract routes swaps to the swapRouter contract and mints/burns/collects to the NFPM. Additionally, it manages all of the NFT liquidity positions (through the ERC721Reciever interface) created by the users in our experiments.
@@ 6 Performance Evaluation
6.1 Experiment Setup
We deploy our system on a computing cluster composed of 8 hypervisors, each running a 12-Core, 130 GiB RAM, VM, connected with 1 Gbps network link. This setup is capable of running around 8000 sidechain miners. Unless stated otherwise, an experiment length is 11 epochs, each of which consists of 30 sidechain rounds (a round lasts 7 sec). Our default meta-block size is 1 MB and a sidechain committee contains 500 miners. We deploy 100 AMM users, generating traffic that arrives at a constant rate of $\rho=\lceil\frac{V_{D}\times b_{t}}{3600\times 24}\rceil$ where $V_{D}=25\times 10^{6}$ is the chosen daily volume of transactions.
In reporting the results, we measure the following metrics:
@@ 6.1 Experiment Setup
    Throughput: Number of transactions processed per second.
@@
    Sidechain transaction latency: The delay between a transaction submission and its appearance in a meta-block. [^14] To obtain an accurate representation of this metric, and thus process a comparable amount of traffic, we empty the transaction queues after the end of each run.
@@
    Mainchain transaction latency: The delay between a transaction submission and its confirmation on Sepolia.
@@
    Payout latency: The delay between a transaction submission and the completion of $\mathsf{TokenBank}$ syncing in the epoch in which this transaction has been published on the sidechain. We measure this metric by reporting the sum of the sidechain transaction latency, the time needed to issue the $\mathsf{Sync}$ call, and the time needed to process the transaction encapsulating the $\mathsf{Sync}$ call on the mainchain.
@@
    Gas cost: The average number of gas units paid to process core transactions.
  6. 6.
@@
6.2 Comparison with the Baseline
We compare $\mathsf{ammBoost}$ against a baseline, which is a deployment of Uniswap V3 on Sepolia, as mentioned earlier.